complete the sequence:
1,(2,3),2,(5,6),4,(11,30),26,(?,?),?.

Answer by fcabanski(1391)

A sequence can have more than one rule, so this is only one possible solution.


The first number in each () is the sum of the numbers in the previous (). For example, 5 = 3+2 and 11 = 6+5.


The second number in () is the product of the numbers in the previous (). 6=2*3 and 30=6*5.


So the next () contains 11+30 and 11*30: (41, 330)


The next number after each () is the difference of the numbers inside the () plus the sum of all the numbers outside the () preceding it.


For example, 2 = 3-2+1 and 4 = 6-5+2+1 and 26 = 30-11+4+2+1


The final number is 330-41+26+4+2+1 = 322.


Solve problems like these by looking for patterns in the way the numbers progress.